Set the Payment Frequency and Compounding Frequency both to 1. Choose an interest rate of 0.10 and an age of 30. Enter payment amount of $1000, $2000, $3000, $4000, and $5000. Note the amount that each investment grows to.
As the payment amount increases by $1000, does the amount that the investment grows to also increase at a uniform rate? Explain.
If the payment amount doubles, what happens to the final amount?
If the payment amount triples, what happens to the final amount?
Summarize the relationship between the payment amount and the final amount of the investment.
1
Expert's answer
2020-05-05T20:45:21-0400


Сalculate according to the formula

"A=S\times(i+\frac{i}{(1+i)^n-1})"


"1000=S\times(0.001+\frac{0.001}{(1+0.001)^{30}-1})"

S=29 540


"2000=S\times(0.001+\frac{0.001}{(1+0.001)^{30}-1})"

S=59 080


"3000=S\times(0.001+\frac{0.001}{(1+0.001)^{30}-1})"

S=88 620


"4000=S\times(0.001+\frac{0.001}{(1+0.001)^{30}-1})"

S=118 160


"5000=S\times(0.001+\frac{0.001}{(1+0.001)^{30}-1})"

S=147 700


Yes, the payment amount is doubled and the total amount is doubled. similarly, and if the payment amount increases by 3 times, then the total amount by times increases and so on, and so on.


It turns out a direct proportional relationship between the payment amount and the total amount